how many liters of oxygen are needed to react when 8.23 x 10^26 molecules of water are produced at stp?


1
Expert's answer
2021-02-26T05:58:40-0500

At STP, one mole of any gas occupies a volume of 22.4 L.

Thus 3.36 L of oxygen (O2) gas correspond to:

Moles of oxygen = 3.36 L × (1 mol / 22.4 L) = 0.15 mol

Moles of oxygen = 0.15 mol


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Oxygen molecules = (0.15 mol of oxygen) × (6.022×1023 molecules / 1 mol) = 9.033×1022 O2 molecules

Oxygen molecules = 9.033×1022 O2 molecules